OpenStudy (anonymous):

complex fraction 1/a-1/b over a/b-b/a I'm not getting the right answer

OpenStudy (anonymous):

You're talking about this, yes?\[\frac{\frac{1}{a}-\frac{1}{b}}{\frac{a}{b}-\frac{b}{a}}\]I'd suggest adding the fractions of the numerator and the fractions of the denominator together as you would normally. See where that takes you. :)

OpenStudy (anonymous):

PS - what answer did you get that you think is wrong?

OpenStudy (anonymous):

In other words, your first step should be the following.\[\frac{\frac{1}{a}\cdot\frac bb-\frac{1}{b}\cdot\frac aa}{\frac{a}{b}\cdot\frac aa-\frac{b}{a}\cdot\frac bb}\]

OpenStudy (anonymous):

I added them together normally and then I did the multiplication. I got b-a/\[a^{2}-b^2\]. That answer is wrong. The one at the back of the text is - 1over a+b

OpenStudy (anonymous):

I did it a few times and I keep getting that same answer. I must be doing it incorrectly.

OpenStudy (anonymous):

You're on the right track...you just need to factor your denominator and simplify it. \(a^2-b^2\) is a difference of two squares.

OpenStudy (anonymous):

thank you so much